OpenStudy (trisarahtops):

The graph of f ′′(x) is continuous and decreasing with an x-intercept at x = 0. Which of the following statements is true? The graph of f has a relative maximum at x = 0. The graph of f has a relative minimum at x = 0. The graph of f has an inflection point at x = 0. The graph of f has an x-intercept at x = 0.

OpenStudy (trisarahtops):

@ganeshie8

OpenStudy (trisarahtops):

@Conqueror

OpenStudy (trisarahtops):

@IrishBoy123

OpenStudy (trisarahtops):

Hi :) is this kind of math familiar to u?

OpenStudy (irishboy123):

none of them are strictly true, IMHO but get a second opinion, i'm crap at maths

OpenStudy (phi):

I know y= -x^3 would give y' = -3x^2 y'' = -6x if we plot y= -6x, it "is continuous and decreasing with an x-intercept at x = 0" and in this case, x=0 is an inflection point.